OpenStudy (anonymous):

What is the x-intercept of the graph: 3x – 9y = 18. A.(0, -2) B.(0, 6) C.(6, 0) D.(-2, 0)

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

3x – 9y = 18 3x – 9(0) = 18 3x - 0 = 18 3x = 18 x = 18/3 x = 6 So the x-intercept is (6,0)

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome